OpenStudy (emogirlatemoooocow):

What is the area of the shaded part of the parallelogram? A. 288 m² B. 432 m² C. 444 m² D. 576 m²

OpenStudy (midhun.madhu1987):

Figure???

OpenStudy (emogirlatemoooocow):



OpenStudy (midhun.madhu1987):

Area of shaded region = Area of Parallelogram - area of Square

OpenStudy (midhun.madhu1987):

Area of a parallelogram = base * height = 24 * 18 = Area of Square = side * side = 12 * 12 =

OpenStudy (midhun.madhu1987):

Hope you can do it from here..

OpenStudy (anonymous):

A?